kaylak:

studying for the SAT. If I disagree with an answer or still don't understand why it is wrong, I will post here. I have 2 questions so far.

kaylak:

@Shadow

kaylak:

If -4<=x<=2 and y=1-x^2, what number is obtained when the smallest possible value of y is subtracted from the largest possible value of y^2?

kaylak:

the answer is 16. I have 15. Can y not=0?

jhonyy9:

1. how many will be the smallest possible value of y ?

gabethebabe:

-4 is less than 0 tho

jhonyy9:

2. how many will be the largest possible value of y^2 ?

jhonyy9:

x=[-4,2] ,yes ?

jhonyy9:

how you get the smallest possible value of y ?

jhonyy9:

@ThisGirlPretty do you understand it now ?

kaylak:

okay do y'all not see the problem above so 1-4^2=15 -4 is the value of x 1-1^2=0 15-0=15 you have to solve the equation for y and subtract both y's but I don't agree with the SAT book being 16 and not 15 because even if you did 0 instead of 1 15-1=14 so idk

jhonyy9:

please answer my above wrote question - how many will be the smalest possible value of y ?

jhonyy9:

for x=2 you get the smallest possible value of y y = 1-2^2 y = 1 -4 y = -3 yes ?

jhonyy9:

2. how many will be the largest possible value of y^2 ?

kaylak:

that's what they are looking for a square root has 2 answers a negative and a positive the negative is for example -2*-2 which is 4 everything in here is as if it was in parentheses every positive number has 2 square roots so although I get what you are trying to get at but it is taught that unlike otherwise stated it is a positive number so that's how I have to do the math to get the right answer and if that is how the sat works, then I must do it that way to get the best score

kaylak:

I understand now. 1-16=-15 1-0^2=1 but my only question is 1-1^2 is 0 wouldn't 0 be lower?

TheSmartOne:

The two end points may not necessarily be the smallest and largest. If it was, that would be too easy of a question. when x=-4 then y=-15 when x=-3 then y=-8 when x=-2 then y=-3 when x=-1 then y=0 when x=0 then y=1 when x=1 then y=0 when x=2 then y=-3 The largest y value is 1, and the smallest y value in our domain is -15 1-(-15) = 1+15 = 16 Does that make sense to you now? You're dealing with a quadratic function, the endpoints they give you may not necessarily be the minimum and maximum or whatever

TheSmartOne:

Also just keep in mind, one question per post (:

kaylak:

I do understand now.

TheSmartOne:

Great! :)